OpenStudy (anonymous):

A population of bacteria is growing according to the function below, where t is in hours. How many hours will it take for the population to grow to 1,000 bacteria? Round your answer to the nearest integer. Do not include units in your answer. B(t) = 50 x e1.5t

OpenStudy (anonymous):

4*e^0.8t = 100 e^0.8t=25 ln (e ^ 0.8t) = ln (25) 0.8t * ln (e) = ln (25) 0.8t * 1 = ln (25) t = ln(25) / (0.8) = 1.25 * ln (25) t= 1.25 * 3.2189 .: t=4.0236